# GVGAI-LLM: Evaluating Large Language Model Agents with Infinite Games

### v1

@@ Introduction
Large language models (LLMs) have shown impressive capabilities in open-ended generation, planning, and reasoning tasks (Achiam et al. 2023; Liu et al. 2024a). Recently, researchers have begun deploying LLMs as interactive agents in sequential decision making environments such as robotics, web navigation, and games (Achiam et al. 2023; Liu et al. 2024a; Yang, Kleinman, and Harteveld 2025). While benchmarks such as GLUE (Wang et al. 2019b), SuperGLUE (Wang et al. 2019a), and MMLU (Hendrycks et al. 2021) evaluate static language understanding and broad domain knowledge, HumanEval (Chen et al. 2021) assesses code generation proficiency, HELM (Liang et al. 2023) examines safety and robustness, and instruction-following frameworks like ALFWorld (Shridhar et al. 2021) and BabyAI (Chevalier-Boisvert et al. 2019) focus on executing textual commands. None of these benchmarks, however, measure decision making competence in structured symbolic domains with game-style logic, real-time decision, and spatial reasoning.
To address this gap, we introduce GVGAI-LLM, a benchmark that adapts the General Video Game AI (GVGAI) framework (Perez-Liebana et al. 2019; Liébana et al. 2019) into a testbed tailored for LLM agents. Unlike existing benchmarks focused on task completion or instruction following, GVGAI-LLM emphasizes goal-directed agent behavior in reactive, rule-based 2D games. The GVGAI framework’s diverse game dynamics and formal Video Game Description Language (VGDL) specifications (Ebner et al. 2013; Schaul 2013) make it ideal for evaluating how well LLMs can parse structured environments, comply with implicit goals, and act.
@@
Games have recently emerged as powerful benchmarks for evaluating a broad range of LLM capabilities, including reasoning, planning, spatial understanding, and multimodal integration. BALROG (Paglieri et al. 2025) assesses LLM agentic behavior across six environments, such as BabyAI (Chevalier-Boisvert et al. 2019) and Nethack (Küttler et al. 2020). Several other benchmarks, SmartPlay (Wu et al. 2024), AgentBench (Liu et al. 2024b), DSGBench (Tang et al. 2025), GTBench (Duan et al. 2024), and LMGame-Bench (Hu et al. 2025), evaluate decision-making and reasoning through a variety of games. GameTraversalBenchmark (Nasir, James, and Togelius 2024) focuses on planning via navigation in procedurally generated levels, while ARC-AGI (Chollet et al. 2024, 2025) emphasizes abstract reasoning through structured puzzles.
Other benchmarks target multimodal understanding, such as Atari-GPT (Waytowich et al. 2024), VideoGameBench (Zhang et al. 2025), and MMBench (Liu et al. 2024c), which require interpreting visual inputs alongside text. These benchmarks highlight the breadth of skills being evaluated in games, yet many rely on either static environments, pre-defined trajectories, or focus primarily on instruction following or multimodal grounding.
@@
The original GVGAI framework (Perez-Liebana et al. 2019; Liébana et al. 2019) provides a unified testbed for evaluating general game-playing agents across diverse environments. It emphasized planning, reactivity, and generalization. Game rules and levels are represented by VGD (Schaul 2013), allowing for the creation of new games and levels to evaluate agents (Khalifa et al. 2016), whereas existing benchmarks typically provide only a single or a fixed set of environments.
GVGAI has been extensively used to evaluate planning agents and reinforcement learning methods, emphasizing generalization, reactivity, and symbolic reasoning (Hu et al. 2022). While it was not originally designed for language-based agents, its modular, rule-based design enables structured interactions that are well-suited for LLM prompting. Modifying GVGAI to accommodate LLM agents enables direct comparison between LLMs, tree search, and RL agents in a shared symbolic space.
@@
VGDL (Ebner et al. 2013; Schaul 2013) is a game description language that encodes game rules and levels/scenes as a compact string of ASCII characters. The Translator module converts VGDL rule specifications into natural language descriptions. It parses object definitions, interaction rules, and termination conditions from the VGDL script. For example, a rule such as avatar > key = > killSprite is rendered as: “If the avatar touches a key, the key disappears and the avatar obtains it.” These textual rule descriptions are provided to downstream modules to support language-based reasoning, as shown in Figure 2.
@@
DeepSeek-chat 16.7% 0.0% 20.0% 0.0% 0.0% 23.3%
Deepseek-r1 0.0% 0.0% 54.5% 0.0% 50.0% 28.6%
@@
In the second experiment, we compare multiple LLMs on a curated subset of six representative games: zelda, aliens, boulderdash, realsokoban, escape, and sokoban with five levels per game. These games were chosen to span a variety of mechanics and difficulty degrees, ranging from real-time action (e.g., aliens, zelda) to spatial puzzles (sokoban, realsokoban) and open-ended navigation tasks (escape). The selection was informed by procedural content generation (PCG) design taxonomies (Shaker, Togelius, and Nelson 2016; Yannakakis and Togelius 2018), ensuring diversity (Li et al. 2025) in spatiality, interactivity, and temporal structure.
@@ Discussion
More fundamentally, we observe that LLMs lack the capacity for path planning in the algorithmic sense. Unlike traditional methods such as A, which search over possible trajectories to compute optimal paths, LLMs operate with limited forward context and no internal simulation of future states. While a reasoning-aware model might theoretically learn to emulate planning through intermediate steps, our outputs do not reveal such structured traces. This limitation aligns with findings from GameTraversalBenchmark (Nasir, James, and Togelius 2024), where LLM agents similarly failed to perform multi-step navigation tasks.

### v2

@@ Introduction
Large language models (LLMs) have shown impressive capabilities in open-ended generation, planning, and reasoning tasks (achiam2023gpt; liu2024deepseek). Recently, researchers have begun deploying LLMs as interactive agents in sequential decision making environments such as robotics, web navigation, and games (achiam2023gpt; liu2024deepseek; 10974629). While benchmarks such as GLUE (wang2018glue), SuperGLUE (NEURIPS2019_4496bf24), and MMLU (hendrycks2021measuring) evaluate static language understanding and broad domain knowledge, HumanEval (chen2021evaluatinglargelanguagemodels) assesses code generation proficiency, HELM (liang2023holistic) examines safety and robustness, and instruction-following frameworks like ALFWorld (shridhar2021alfworld) and BabyAI (chevalier-boisvert2018babyai) focus on executing textual commands. None of these benchmarks, however, measure decision making competence in structured symbolic domains with game-style logic, real-time decision, and spatial reasoning.
To address this gap, we introduce GVGAI-LLM, a benchmark that adapts the General Video Game AI (GVGAI) framework (perez2019general; liebana2019general2) into a testbed tailored for LLM agents. Unlike existing benchmarks focused on task completion or instruction following, GVGAI-LLM emphasizes goal-directed agent behavior in reactive, rule-based 2D games. The GVGAI framework’s diverse game dynamics and formal Video Game Description Language (VGDL) specifications (ebner2013towards; schaul2013video) make it ideal for evaluating how well LLMs can parse structured environments, comply with implicit goals, and act.
@@
Games have recently emerged as powerful benchmarks for evaluating a broad range of LLM capabilities, including reasoning, planning, spatial understanding, and multimodal integration. BALROG (paglieri2025balrog) assesses LLM agentic behavior across six environments, such as BabyAI (chevalier-boisvert2018babyai) and Nethack (kuttler2020nethack). Several other benchmarks, SmartPlay (wu2024smartplay), AgentBench (liu2024agentbench), DSGBench (tang2025dsgbench), GTBench (duan2024gtbench), and LMGame-Bench (hu2025lmgame), evaluate decision-making and reasoning through a variety of games. GameTraversalBenchmark (nasir2024gametraversalbenchmark) focuses on planning via navigation in procedurally generated levels, while ARC-AGI (chollet2024arc; chollet2025arc) emphasizes abstract reasoning through structured puzzles.
Other benchmarks target multimodal understanding, such as Atari-GPT (waytowich2024atari), VideoGameBench (zhang2025videogamebench), and MMBench (liu2024mmbench), which require interpreting visual inputs alongside text. These benchmarks highlight the breadth of skills being evaluated in games, yet many rely on either static environments, pre-defined trajectories, or focus primarily on instruction following or multimodal grounding.
@@
The original GVGAI framework (perez2019general; liebana2019general2) provides a unified testbed for evaluating general game-playing agents across diverse environments. It emphasized planning, reactivity, and generalization. Game rules and levels are represented by VGD (schaul2013video), allowing for the creation of new games and levels to evaluate agents (khalifa2016general), whereas existing benchmarks typically provide only a single or a fixed set of environments.
GVGAI has been extensively used to evaluate planning agents and reinforcement learning methods, emphasizing generalization, reactivity, and symbolic reasoning (hu2022reinforcement). While it was not originally designed for language-based agents, its modular, rule-based design enables structured interactions that are well-suited for LLM prompting. Modifying GVGAI to accommodate LLM agents enables direct comparison between LLMs, tree search, and RL agents in a shared symbolic space.
@@
VGDL (ebner2013towards; schaul2013video) is a game description language that encodes game rules and levels/scenes as a compact string of ASCII characters. The Translator module converts VGDL rule specifications into natural language descriptions. It parses object definitions, interaction rules, and termination conditions from the VGDL script. For example, a rule such as avatar > key = > killSprite is rendered as: “If the avatar touches a key, the key disappears and the avatar obtains it.” These textual rule descriptions are provided to downstream modules to support language-based reasoning, as shown in Figure 2.
@@
Deepseek-chat 16.0% 0.0% 20.0% 0.0% 0.0% 24.0%
Deepseek-r1 0.0% 0.0% 24.0% 0.0% 12.0% 16.0%
@@
In the second experiment, we compare multiple LLMs on a curated subset of six representative games: zelda, aliens, boulderdash, realsokoban, escape, and sokoban with five levels per game. These games were chosen to span a variety of mechanics and difficulty degrees, ranging from real-time action (e.g., aliens, zelda) to spatial puzzles (sokoban, realsokoban) and open-ended navigation tasks (escape). The selection was informed by procedural content generation (PCG) design taxonomies (shaker2016procedural; Yannakakis2018Artificial), ensuring diversity (li2025measuring) in spatiality, interactivity, and temporal structure.
@@ Discussion
More fundamentally, we observe that LLMs lack the capacity for path planning in the algorithmic sense. Unlike traditional methods such as A, which search over possible trajectories to compute optimal paths, LLMs operate with limited forward context and no internal simulation of future states. While a reasoning-aware model might theoretically learn to emulate planning through intermediate steps, our outputs do not reveal such structured traces. This limitation aligns with findings from GameTraversalBenchmark (nasir2024gametraversalbenchmark), where LLM agents similarly failed to perform multi-step navigation tasks.